HYDROGEN BONDING IN CHIRAL MOLECULAR COMPLEX: DIRECT DETECTION OF PROPYLENE OXIDE-WATER ADDUCT IN GAS PHASE

\underline{ZHENG SU} AND YUNJIE XU, Department of Chemistry, University of Alberta, Edmonton, AB, T6G 2G2, Canada.

High resolution spectroscopy and ab initio calculations of chiral molecules and their molecular complexes promise to provide rich structural and dynamical information about this class of molecular systems. In this talk, we will report the detailed rotational spectroscopic and high level ab initio studies of the hydrogen bonded propylene oxide-water (C_{3}H_{6}O-H_{2}O) complex. Two distinct structural conformers of the 1:1 molecular adduct were detected experimentally. Rotational spectra of the four isotopomeric species, namely C_{3}H_{6}O-H_{2}O, C_{3}H_{6}O-DOH, C_{3}H_{6}O-HOD and C_{3}H_{6}O-D_{2}O were recorded for the two conformers. The effect of the methyl group on the stability of the O_{epoxy}^{...}H-O type hydrogen bond will be examined by comparing with the closely related ethylene-water adduct.^{a}

aW. Caminati, P. Moreschini, I. Rossi, and P. G. Favero J. Am. Chem. Soc. 1998, 120, 11144-11148.